OpenStudy (lilday12):

What is the factorization of the trinomial below? -2x3 - 10x2 - 12x A. -2x(x + 3)(x - 2) B. -2x(x - 3)(x + 2) C. -2x(x + 3)(x + 2) D. 2x(x2 + 3)(x + 2)

OpenStudy (lilday12):

Yes

OpenStudy (er.mohd.amir):

-2x(x*x+5x+6)

OpenStudy (er.mohd.amir):

now x*x+3x+2x+6=x(x+3)+2(x+3)=(x+3)(x+2)

OpenStudy (er.mohd.amir):

-2x(x+3)(x+2)

OpenStudy (lilday12):

The expression below is the factorization of what trinomial? Enter the trinomial in descending order. Enter exponents using the caret ( ^ ). For example, you would enter x2 as x^2. -1(x + 5)(x + 6)